For the given matrix A, find a basis for the corresponding eigenspace for the given eigenvalue. 1 6 6 6 1 - 6, A= 7 6 6 13 1)How is option D the correct answer? I perform R2-R3 = R3 which results in all zeros for R3 Then I perform R1-R2 = R2 which results in all zeros for R2. Resulting in (1 -1 -1) in R1 and I get vectors (1,0,1) (1,1,0) but keep getting wrong result. Any help is appreciated
Answer #1

find eigenspact, [A-XIJX = 0 To ue solve system for a=7 -6 ie 9.- (R2 R2+R, R3+ R3-R1) 6. -6 6. 6. 24 2) - 671+ 672 t 63=0 -x